jhonyy9 (jhonyy9):

,,1/2 can being a part of every real numbers " true or false ?

OpenStudy (experimentx):

for example??

jhonyy9 (jhonyy9):

1= 2*(1/2) =2/2 =1 2= 4*(1/2) =4/2 =2 3= 6*(1/2) =6/2 =3 . . . - so i dont know if in this way i have thought it right ???

OpenStudy (experimentx):

most likely not ... real numbers include irrational and transcendental numbers.

OpenStudy (experimentx):

pi = 2pi*(1/2) <--- probably you will get something like this ... I don't know if this works for you.

jhonyy9 (jhonyy9):

so in this way what is your opinion - please ! - can being writing every real numbers using 1/2 like a part of thise ?

OpenStudy (experimentx):

in general you can always have a = (2a)*1/2 what is the criteria for (2a)?

jhonyy9 (jhonyy9):

have not nothing criteria for (2a) so than how do you think please ,can being this always true ?

OpenStudy (experimentx):

well ... i would say ... every number is half of it's double. But I'm not quite suitable for this topic.

jhonyy9 (jhonyy9):

ok - thank you very much ! - so the criteria for (2a) what we know is that need to be an indifferent real number --- so in this case ,what is your opinion ,please ?

OpenStudy (experimentx):

i would say yes!! but don't rely on me ... I am not suitable for these topic.

jhonyy9 (jhonyy9):

thank you ! ... so hope so much that i will get again opinions from this and ,so will can knowing how think this more persons on this site ... !!! ???

OpenStudy (experimentx):

yup!!

OpenStudy (anonymous):

Hello jhon - when they say "part of" they mean ADDITIO and SUBSTRACTION, the answer is YES

OpenStudy (anonymous):

ANY NUMBER X: X = 1/2 + (X -1/2) That's all

OpenStudy (anonymous):

hee-looo -oo anybody heere ?...

jhonyy9 (jhonyy9):

some different opinion ??? ---please ???

jhonyy9 (jhonyy9):

,,Mikael" i think that you have wrote it right ,but can you prove it too EASY ???

jhonyy9 (jhonyy9):

...or not is necessary

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

this debate is retarded

OpenStudy (anonymous):

This guy is A) Champion B) Troll C) Thinks that any of these two things make him anything but a piece of organic matter+warm liquid which he clearly is

Directrix (directrix):

The set of Real numbers is closed with respect to the operation of multiplication. I think that it is true that every Real number can be written as a product of 1/2 and another Real number. That other Real number may seem contrived. ---- Example: Square Root (5) = (1/2) * 2[ Square Root (5) ]

OpenStudy (anonymous):

true

OpenStudy (anonymous):

true.. real number including.. 1)whole number (1,2,3,...) 2) rational number( 1/2,3/4) 3)irrational number( \[\sqrt{3}\]

OpenStudy (aravindg):

yes 1/2 can be

OpenStudy (aravindg):

its just 1-1 matching you see for every real no:x , 2x also exists